OpenStudy (anonymous):

a sock drawer contains 10 white sock, 6 black and 4 brown mismatched socks cooper randomly picks 2 socks without replacing the first. what is the probability that he gets a white pair

OpenStudy (anonymous):

first one is white AND second one is white GIVEN that first one is white

OpenStudy (anonymous):

so \[\frac{10}{20}\times \frac{9}{19}=\frac{9}{38}\]

OpenStudy (anonymous):

THX!